### v1

@@ 4.2 Reduced Use of Weight Regularization
As discussed in Section 3.3, weight regularization reduces PANNs’ accuracy. One possible solution is to use minimal weight regularization combined with early stopping. However, it will lead to poor generalization, which can reduce the accuracy of backbone models, thereby limiting PANN’s accuracy by reducing its upper bound.
Therefore, we choose Mixup as a remedy method. Mixup Zhang et al. (2018b) introduces globally linear behavior in-between data manifolds. It can reduce unexpected behavior of neural networks and enhance generalization and robustness by expanding the input set. Our experiments show that Mixup can somewhat compensate for the reduced accuracy due to poor regularization. The test results based on minimal regularization and Mixup are presented in Section 5.2, which shows good performance on some model structures and precision. However, it is important to note that this approach is only effective when regularization is not crucial. For deeper models, the problem of low backbone accuracy under poor regularization is still challenging. Besides, Mixup may decrease PANN’s accuracy for certain models, such as Shufflenetv2. Thus, we regard use of low weight regularization and Mixup as a trade-off.
@@ 5 Experiments
We conduct experiments to demonstrate effects of the perturbations on irrelevant information in the input background as well as the effectiveness of our solutions to improve sturdiness and the accuracy of low-precision PANN. The main results are shown in this section. Additional experiment results can be found in Appendix A.
5.1 Effects of Approximation Errors illustrated by PGD attacks
(a) Success rate of PGD attacks on PANN
(b) Percentage of $\delta$ only affects PANN
Figure 4: (a) PGD attacks on both regions can have high success rates on PANN. (b)PGD attacks on large-gradient regions have high success rates in backbone models, but attacks on small-gradient regions can hardly affect the backbone model.
Section 3.2 mentioned that approximation errors lead to intra-model perturbations on irrelevant information in the input background and are extremely harmful to PANN. To prove that, we create adversarial samples with perturbation added to irrelevant information in the input background and information contributing to the outputs separately. “Irrelevant” means the information is filtered and has little effect on model outputs and corresponds to image background in images. If we create adversarial samples with perturbation added to input backgrounds, the perturbations will be transferred throughout PANN by approximation errors and enlarge the error by multiplication. This will severely impact PANN but have little effect on the backbone models. However, if we attack objects in images (corresponding to information contributing to the outputs), both PANN and backbone will be affected.
In our experiment, we create adversarial samples with perturbation added to objects and backgrounds separately by preserving part of input perturbations $\delta$ according to magnitude of gradients in the backbone model: Attacks on large-gradient regions preserve perturbations on pixels with the top 20% gradient magnitude (usually the object); Attacks against small-gradient regions preserve pixels with the minimum 30% gradient magnitude (usually the image background). For each $\alpha$, we randomly select 5000 images.
Figure 4 shows that attacks on backgrounds and objects all have a high success rate on PANN. However, only attacks on objects can easily affect backbone models. When the step length is very small, attacks on both large-gradient and small-gradient regions will first only mislead PANN. Immediately after the step length increases, attacks on large-gradient regions also affect the backbone model. In contrast, attacks on small-gradient regions only have 32% success rate on backbone models even though the step length is 40 times the beginning. This result is a good proof of the assumption in Section 3.2. It demonstrates that PANN suffers from approximation errors in the irrelevant information in the input background, opening the possibility of dealing with approximation errors by leveraging this information.
5.2 Accuracy Evaluation
Table 2: Accuracy of PANN on ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2 trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (bb) and PANN with precision $2^{-\beta}$ ($2^{-8}$ and $2^{-9}$). Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
Vanilla Mixup NGNV Mixup+NGNV
$\beta$ wd 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4 0 1e-4 5e-4
[…]
bb 90.39 91.57 92.14 91.30 91.60 92.17 91.10 91.70 92.49 91.49 91.59 92.16
[…]
Table 3: Accuracy of PANN with precision $2^{-\beta}$ and the backbone (bb) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18).
@@ 5.2 Accuracy Evaluation
Evaluation Setup: We conduct tests on PANN with ResNet He et al. (2015), DLA Yu et al. (2018), MobilenetV2 Sandler et al. (2019), and ShufflenetV2 Ma et al. (2018) models for classification tasks on CIFAR-10. We also test ResNet on CIFAR-100 and Tiny Imagenet dataset Krizhevsky et al. (2009); Le & Yang (2015). Training on CIFAR dataset takes 200 epochs, and on Tiny Imagenet takes 150 epochs. All learning rates begin with 0.1, and were multiplied by 0.1 on epochs 100 and 150 on CIFAR dataset, multiplied by 0.1 on epochs 50 and 100 on Tiny Imagenet. The momentum is 0.9. We test weight decay on 0, 0.0001, and 0.0005. Note that the ResNet-18 models removed the Maxpool layer to suit small images. $\lambda$ for Mixup are selected from distribution $\operatorname{Beta}(0.5,0.5)$. The approximation intervals are $[-50,50]$ for CIFAR and $[-100,100]$ for Tiny Imagenet, except the case values exceed the approximation interval. The average accuracy of the backbone models and PANN with precision $2^{-8}$ and $2^{-9}$ on CIFAR-10 dataset are presented in Table 2. The experiments on CIFAR-100 (C100) and Tiny Imagenet are shown in Table 3. The highest accuracy for each model and approximation precision are highlighted in bold.
@@
Effects of Weight Regularization: From Table 2 and Table 3, we observe that weight regularization can severely destroy the model’s resistance to approximation errors in various models and precisions. For example, PANN on models trained with weight decay 0 outperforms those with weight decay 5e-4 in nearly all vanilla models, suggesting minimal weight decay. Note that this discipline doesn’t always hold. Poor regularization caused by no regularization can limit backbone models’ accuracy, thus reducing PANN’s accuracy by limiting the upper bound. For example, ResNet-20 models with weight decay 1e-4 can outperform those with zero weight decay in precision $2^{-9}$.
@@
Evaluation of Minimal Regularization with Mixup: Table 2 and Table 3 shows that Mixup can effectively compensate for the lack of generalization and reduced backbone accuracy caused by poor regularization. Therefore, models can benefit from strong resistance to approximation errors which come with minimal weight regularization. This method can improve PANN performance in many cases. For example, Mixup can outperform vanilla models on precision $2^{-9}$ with zero weight decay in ResNet-20 and DLA-34. However, this advantage may not always hold. In some cases, Mixup might be harmful to PANN (eg: DLA-34, weight decay 0, precision $2^{-8}$).
@@
Evaluation of AT-like Solution: Table 2 and Table 3 prove that NGNV (or NGNV with Mixup), can achieve the best accuracy in all precisions (bolded). It can also improve PANN accuracy significantly in almost all cases, even if the weight decay is large. For instance, on Shufflenetv2 and Mobilenetv2, our models can improve PANN accuracy by up to 60% on weight decay 0 and precision $2^{-8}$, by over 40% on weight decay 1e-4 and precision $2^{-9}$.
@@
In a word, by combining these methods, models can exhibit strong “sturdiness” to approximation errors. We can significantly improve low-precision PANN’s accuracy and thus reduce the time cost. Table 1 illustrates some time costs for PANN with different precision on different models. Previous schemes require at least $2^{-12}$ precision to achieve satisfactory accuracy, but our method only needs $2^{-9}$ (Kindly refer to the appendix for the selection of parameters). The time required for PANN with our models to achieve the state-of-the-art Lee et al. (2021a; 2022b) accuracy is given in Figure 1. Our solution can reduce the PANN time cost by 40% to 60% compared to models provided by these schemes. This improvement makes it easier to design cryptographic schemes for various scenarios.
@@ 6 Conclusions
We uncovered crucial insights into the nature of “sturdiness”, a notion we introduced as a network’s resistance to approximation errors. We identified the intricate relationship between “sturdiness” and adversarial robustness and introduce an AT-like method tailored to enhance sturdiness. Additionally, we discovered that weight regularization negatively affect “sturdiness”, and explored alternative generalization strategies.
@@
We conclude with a few open problems that warrant deeper exploration, namely, establishing a precise mathematical relationship between model accuracy and approximation error, understanding the implications of weight regularization, and design of solutions applicable at different precision levels.
@@ Appendix A Additional Experiment Results
Table 4: Accuracy of PANN on CIFAR-10 dataset (ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2) trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (bb) and PANN with precision $2^{-\beta}$. Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
@@
This section gives more results of the accuracy of PANN on models trained with different methods and weight decays. The experiment results on CIFAR-10 dataset are shown in Table 4, on CIFAR-100 (C100) and TinyImagenet is shown in Table 6. Note that PANN’s accuracy is very close to the backbone models for high precision like $2^{-11}$ and $2^{-12}$, which reduce the advantage of small weight regularization.
Table 5: Accuracy and time cost of PANN with precision $2^{-\beta}$ and the backbone (bb) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18). The results with the highest accuracy for each model and approximation precision are bolded.
@@
We also give parameters $r$ and $\lambda$ for our NGNV in Table 6 to help the reader reproduce our results.
Table 6: Best parameters ($r$, $\lambda$) for NGNV on different weight decay and methods
@@
The accuracy and time cost for models released by us and the state-of-the-art schemes on different precisions are presented in Table 7. (Time cost for ResNet-20 has been given in Table 1.)
Table 7: Comparation of Models Provided by us and Models offered by the State-of-the-Art Scheme Lee et al. (2022b) (Classification Accuracy on CIFAR-10)

### v2

@@ 5.2 Reduced Use of Weight Regularization
As discussed in Theorem 2.2, weight regularization increases PANNs’ test loss and reduces accuracy as training processes. One potential solution is to use minimal weight regularization combined with early stopping. However, it will lead to poor generalization, which can reduce the accuracy of backbone models (baseline) and ultimately limit PANN’s accuracy by decreasing its upper bound.
To address this, we choose Mixup as a remedy method. Mixup [30] introduces globally linear behavior in-between data manifolds. It can reduce unexpected behavior of neural networks and enhance generalization and robustness by expanding the input set. Our experiments show that Mixup can somewhat compensate for the reduced accuracy due to poor regularization. The test results based on minimal regularization and Mixup are presented in Section 5.2, which shows good performance on some model structures and precision. However, it is important to note that this approach is only effective when regularization is not crucial. For deeper models, the problem of low backbone accuracy under poor regularization is still challenging. Besides, Mixup may decrease PANN’s accuracy for certain models, such as Shufflenetv2. Thus, we regard use of low weight regularization and Mixup as a trade-off.
@@ 6 Experiments
We conduct experiments to demonstrate the effectiveness of our solutions in improving sturdiness and the performance of PANN. The main results are shown in this section and additional experiment results can be found in Appendix A.
6.1 Test on Normal Structure without Fine-tuning
TABLE III: Accuracy of PANN with error bound $2^{-\beta}$, and the backbone (baseline) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18).
@@ 6.1 Test on Normal Structure without Fine-tuning
Evaluation Setup: We conduct tests on PANN with ResNet [32], DLA [33], MobilenetV2 [34], and ShufflenetV2 [35] models for classification tasks on CIFAR-10. We also test ResNet on CIFAR-100 and Tiny Imagenet dataset [36, 37]. Training on CIFAR dataset takes 200 epochs, and on Tiny Imagenet takes 150 epochs. All learning rates begin with 0.1, and were multiplied by 0.1 on epochs 100 and 150 on CIFAR dataset, multiplied by 0.1 on epochs 50 and 100 on Tiny Imagenet. The momentum is 0.9. We test weight decay on 0, 0.0001, and 0.0005. Note that the ResNet-18 models removed the Maxpool layer to suit small images. $\lambda$ for Mixup are selected from distribution $\operatorname{Beta}(0.5,0.5)$. The approximation intervals are $[-50,50]$ for CIFAR and $[-100,100]$ for Tiny Imagenet, except the case values exceed the approximation interval. The average accuracy of the backbone models and PANN with precision $2^{-8}$ and $2^{-9}$ on CIFAR-10 dataset are presented in Table II. The experiments on CIFAR-100 (C100) and Tiny Imagenet are shown in Table III. The highest accuracy for each model and approximation precision are highlighted in bold.
@@
Effects of Weight Regularization: From Table II and Table III, we observe that weight regularization can severely destroy the model’s resistance to approximation errors in various models and precisions. For example, PANN on models trained with weight decay 0 outperforms those with weight decay 5e-4 in nearly all vanilla models, suggesting minimal weight decay. Note that this discipline doesn’t always hold. Poor regularization caused by no regularization can limit backbone models’ accuracy, thus reducing PANN’s accuracy by limiting the upper bound. For example, ResNet-20 models with weight decay 1e-4 can outperform those with zero weight decay in precision $2^{-9}$.
@@
Evaluation of Minimal Regularization with Mixup: Table II and Table III shows that Mixup can effectively compensate for the lack of generalization and reduced backbone accuracy caused by poor regularization. Therefore, models can benefit from strong resistance to approximation errors which come with minimal weight regularization. This method can improve PANN performance in many cases. For example, Mixup can outperform vanilla models on precision $2^{-9}$ with zero weight decay in ResNet-20 and DLA-34. However, this advantage may not always hold. In some cases, Mixup might be harmful to PANN (e.g.: DLA-34, weight decay 0, precision $2^{-8}$).
@@
Evaluation of Negative Inputs Leakage Solution: Table II and Table III prove that NGNV (or NGNV with Mixup), can achieve the best accuracy in all precisions (bolded). It can also improve PANN accuracy significantly in almost all cases, even if the weight decay is large. For instance, on Shufflenetv2 and Mobilenetv2, our models can improve PANN accuracy by up to 60% on weight decay 0 and precision $2^{-8}$, by over 40% on weight decay 1e-4 and precision $2^{-9}$.
Cross-work comparison: Table IV and V provide a cross-work comparison for different schemes. Note that due to the large amount of fine-tuning and structural adjustment involved in the different schemes, it’s difficult to assess the magnitude of errors in them. So we only present the highest accuracy obtained. And in this experiment, our scheme uses the same approximation precision as MPCNN [5]. The results indicate that though our models are trained on the original backbone structure without fine-tuning with approximation functions, they can achieve nearly the same accuracy as the baseline when loaded in PANN. In contrast, model modification and fine-tuning are necessary for other high-accuracy schemes (e.g., reducing the number of layers containing activation functions).  [^2] Note that our method has the potential to be integrated with these fine-tuning schemes and enhance their performance, but due to the limited open-source material available in this specific area, we have only conducted tests using a subset of the available schemes.
TABLE IV: Best accuracy of PANN on CIFAR-10 dataset. “Fine-tuned” means if the model is fine-tuned on the PANN. “Self-defined” means the model structure is designed for PANN specifically and cannot be identified as common-used structures.
Model Schemes Fine-tuned Accuracy Baseline
ResNet20 AutoFHE [11] Yes 92.66 92.89
MPCNN [5] No 91.39 91.52
Ours No 92.05 92.02
ResNet56 AutoFHE [11] Yes 93.27 93.49
MPCNN [5] No 93.12 93.26
Ours No 93.77 93.81
ResNet56-Modified LSPCNN [3] Yes 93.27 97.80
Self-defined HCNN [7] 77.55
nGraph-HE [8] 62.10
TABLE V: Best accuracy of PANN on CIFAR-100 dataset
Model Schemes Fine-tuned Accuracy Baseline
ResNet32 AutoFHE [11] Yes 70.81 71.34
MPCNN [5] No 69.50 69.43
Ours No 70.72 70.96
Figure 4: Time cost for PANN with our models and the non-fine-tuning SOTA [15, 5] (to achieve the SOTA accuracy on CIFAR-10).
@@
In a word, by combining these methods, models can exhibit strong “sturdiness” to approximation errors. We can significantly improve low-precision PANN’s accuracy and thus reduce the time cost. Table I illustrates some time costs for PANN with different precision on different models. Previous schemes require at least $2^{-12}$ precision to achieve satisfactory accuracy, but our method only needs $2^{-9}$ (Kindly refer to the appendix for the selection of parameters). The time required for PANN with our models to achieve the state-of-the-art [15, 5] accuracy is given in Figure 4. Our solution can reduce the PANN time cost by 40% to 60% compared to models provided by these schemes. This improvement makes it easier to design cryptographic schemes for various scenarios.
6.2 Test on ReLU Replacement Schemes
Our method is based on standard structures without fine-tuning (mentioned in Section 1.1), which ensures our model can benefits other schemes by improving sturdiness. For instance, our models can serve as pre-trained and teacher models (for distillation) to enhance the accuracy of ReLU Replacement schemes SNL [17] and AutoReP [2].
TABLE VI: The accuracy (%) of SNL, AutoReP, and the two schemes with our pre-trained model. (CIFAR-100, ResNet-18, baseline: 77.9% for ours, 77.8% for SNL and AutoReP)
ReLUs(K) Accuracy(%)
Sphynx [17] 51.2 69.57
25.6 66.13
DeepReduce [38] 49.2 69.5
12.3 64.97
SNL [17] 49.9 73.75
24.9 70.05
15.0 67.17
SNL+Ours 49.9 74.46
24.9 71.18
15.0 68.62
AutoReP [2] 50.0 75.48
12.9 74.92
6.0 73.79
AutoReP+Ours 50.0 78.27
12.9 77.91
6.0 77.81
Table VI presents the performance of AutoReP and SNL with (or without) our pre-trained ResNet-18 model on CIFAR-100 dataset. We also give the accuracy of Sphynx [17], DeepReduce [38] for comparison. Our models are trained on the backbone structure with weight decay 1e-4 for 200 epochs. Learning rates begin with 0.1, multiplied by 0.1 on epochs 100 and 150. The table shows that our models can enhance both the accuracy of both SNL and AutoRep. For example, we can improve the accuracy of AutoRep with 50K, 12.9K, and 6K ReLUs to near the upper bound (baseline).
TABLE VII: The accuracy (%) of ReLU replacement schemes SNL, AutoReP with models pre-trained and fine-tuned with different weight decay (wd) (CIFAR-100, ResNet-18)
wd 0 1e-4 5e-4 ReLUs(K)
SNL 70.6 73.3 70.38 49.90
68.48 71.17 68.33 24.90
AutoRep 72.95 73.44 75.27 50.00
69.43 73.32 73.97 12.90
baseline 71.81 74.5 77.66
We also give the accuracy of SNL and AutoReP with models pre-trained and fine-tuned with different weight decay in Table VII. It shows that approximations with 1-degree (in SNL) and 2-degree (in AutoRep) polynomials also comply with Theorem 2.2, demonstrating that weight decay enlarges the loss increment due to approximation errors. Note that this improvement is achieved without using our training methods in fine-tuning, which means it can be further improved.
6.3 Test on Fixed-Point Truncation-based ReLU
TABLE VIII: The accuracy (%) of fix-point truncation-based ReLU computation (CIFAR-100, ResNet-18)
Vanilla Ours
wd 1e-4 5e-4 1e-4 5e-4
baseline 74.80 77.82 78.63 78.85
lx=6 74.39 73.09 78.33 75.98
lx=8 74.64 76.58 78.38 78.22
lx=10 74.74 77.57 78.38 78.71
lx=12 74.77 77.68 78.57 78.83
lx=14 74.77 77.80 78.64 78.88
lx=16 74.80 77.80 78.65 78.84
The fixed-point bits ($l_{x}$) of ReLU’s inputs are positively correlated with communication overhead in certain MPC schemes due to bit-wise operations such as truncation. By choosing a small $lx$, MPC can significantly reduce communication costs at the expense of accuracy. Our scheme’s increasing model sturdiness can effectively mitigate the accuracy reduction resulting from a small $l_{x}$.
Table VIII shows the accuracy on models with different ReLU inputs bits. The model training uses the same setting as Section 6.2. And we apply the computation methods in [18]. We set the first half of the bits in lx represent integers and the second half represent decimals.
The table shows that this case still experiencing external losses due to weight decay. Though the large weight decay outperforms the small weight decay in some precision because of the baseline difference, the reduced accuracy in weight decay 1e-4 is still smaller than that in 5e-4. Furthermore, by applying our methods in training backbone, models can achieve the desired accuracy at much lower precision, ultimately reducing communication overhead.
@@ 7 Conclusions
We uncovered crucial insights into the nature of “sturdiness”, a notion we introduced as a network’s resistance to approximation errors. We proposed the theorems regarding how approximation errors increase the test loss on PANN and give constraints on the relevant boundaries.
Based on these theorems, we have explored the factors that will affect the model “sturdiness” and drawn two interesting conclusions. First, we found that approximations on ReLU’s negative inputs result in more testing loss compared to approximations on positive inputs, based on which we developed NGNV solution that can enhance “sturdiness” with minimal effect on the backbone model’s performance Second, we discovered that weight regularization negatively affects “sturdiness”, suggesting applying minimal weight regularization and exploring alternative generalization strategies. By combining the methods above, we can significantly improve low-precision PANN’s accuracy, thus reducing time costs.
@@
Beyond these contributions, we also leave some open questions for subsequent research. Firstly, our theorems only target small errors (near zero). Due to the complex accumulation and amplification of noise in the neural networks, approximation errors in some precision (i.e., $\epsilon\leq 2^{-8}$) will lead to results inconsistent with our theorem. Finding a more general mathematical model to address this issue and help with analysis is important. Secondly, similar to AT [39, 40], NGNV suffers from a form of “overfitting”. A model may perform better beyond some precision but have worse accuracy on lower precision (i.e. models trained with small noise may perform poorly on approximation with extremely large errors). How to make models perform well in every accuracy still needs further research.
Overall, this research has opened up a new field and made a great contribution to PPML by providing powerful tools and new directions for the future development of PPML.
@@ Appendix A Additional Experiment Results
TABLE IX: Accuracy of PANN on CIFAR-10 dataset (ResNet-20, Shufflenetv2, DLA-34, Mobilenetv2) trained with weight decay (wd) 0, 1e-4 and 5e-4. We give Top-1 accuracy of the backbone (bb) and PANN with precision $2^{-\beta}$. Training methods include Vanilla, Mixup, and robustness-based solution NGNV. The results with the highest accuracy for each model and approximation precision are bolded.
@@
This section gives more results of the accuracy of PANN on models trained with different methods and weight decays. The experiment results on CIFAR-10 dataset are shown in Table IX, on CIFAR-100 (C100) and TinyImagenet is shown in Table X. Note that PANN’s accuracy is very close to the backbone models for high precision like $2^{-11}$ and $2^{-12}$, which reduce the advantage of small weight regularization.
TABLE X: Accuracy and time cost of PANN with precision $2^{-\beta}$ and the backbone (bb) on CIFAR-100 (ResNet-32) and TinyImagenet (ResNet-18). The results with the highest accuracy for each model and approximation precision are bolded.
@@
We also give parameters $r$ and $\lambda$ for our NGNV in Table XI to help the reader reproduce our results.
TABLE XI: Best parameters ($r$, $\lambda$) for NGNV on different weight decay and methods
@@
The accuracy and time cost for models released by us and the state-of-the-art schemes on different precisions are presented in Table XII. (Time cost for ResNet-20 has been given in Table I.)
TABLE XII: Comparation of PANN [5] with us models and the state-of-the-art non-fine-tune scheme (Classification Accuracy on CIFAR-10)